MOA-II Obsevation Log Monthly Summary 2007/09
Return to index
Previous month
Next month
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 10247-11140 | 894 | 3189 | 184615 |
LMC | 1693-1801 | 109 | 587 | 176100 |
SMC | 433-444 | 12 | 22 | 6600 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 1015 | 3798 | 367315 (102h01m55s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
141 | 156 | 127 | 93 | 354 | 78 | 103 | 285 | 416 | 81 | 126 | 91 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
83 | 82 | 342 | 113 | 82 | 96 | 92 | 75 | 78 | 95 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
27 | 25 | 22 | 18 | 17 | 29 | 105 | 101 | 98 | 29 | 29 | 29 |
Field13 | Field14 | Field15 | Field16 |
29 | 29 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
11 | 11 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAME
CCD TEMPERATURE
Previous month
Next month

VACUUM
Previous month
Next month

TELESCOPE TEMPERATURE
Previous month
Next month

COMMENTS
2007/09/1: DARK D2663(15sec) for FLAT and D2665(60sec), D2666(300sec) are OK / night time = 39840(sec)
2007/09/2: No observation / night time = 39660(sec)
2007/09/3: DARK D2667,D2668(60sec) are OK / night time = 39420(sec)
2007/09/4: No observation / night time = 39300(sec)
2007/09/5: anomaly MOA379(gb15) observation, DARK D2669(60sec) and D2674(60sec), D2670(15sec), D2671(30sec), D2672(45sec), D2673(300sec) are OK / night time = 39120(sec)
2007/09/6: DARK D2677(60sec), D2678(300sec) are OK / night time = 38940(sec)
2007/09/7: DARK D2679(60sec) and D2680(300sec) are OK / night time = 38760(sec)
2007/09/8: DARK D2681(60sec) and D2682(300sec) are OK / night time = 38580(sec)
2007/09/9: DARK D2684(60sec) is OK, take FLAT with many rotator angle (F1926-F1934, F1937-F1963) and for gain (F1967-F1982), take DARK for FLAT (D2685-D2700) / night time = 38400(sec)
2007/09/10: DARK D2701(60sec) and D2702(300sec) are OK / night time = 38280(sec)
2007/09/11: DARK D2703(60sec) is OK / night time = 38040(sec)
2007/09/12: DARK D2704(60sec) and D2705(300sec) are OK, Pumping after observation / night time = 37860(sec)
2007/09/13: No observation, Pumping is going on / night time = 37680(sec)
2007/09/14: DARK D2709(60sec) and D2710(300sec) are OK / night time = 37500(sec)
2007/09/15: DARK D2712(60sec) and D2713-300-a,D2714-300-a,D2717-300-a(300sec) are OK / night time = 37320(sec)
2007/09/16: DARK D2718(60sec) and D2719(300sec) are OK / night time = 37080(sec)
2007/09/17: Pumping before observation, DARK D2720(60sec) and D2721(300sec) are OK / night time = 36900(sec)
2007/09/18: DARK D2722(60sec) and D2723(300sec) are OK / night time = 36720(sec)
2007/09/19: DARK D2725(60sec) and D2726(300sec) are OK, GB is close to MOON / night time = 36540(sec)
2007/09/20: DARK D2727(15sec) and D2728(30sec), D2729(60sec), D2730(300sec) are OK, Pumping before observation, GB is close to MOON / night time = 36300(sec)
2007/09/21: Pumping before observation / night time = 36120(sec)
2007/09/22: change new vacuum sensor, Pumping on going, No observation / night time = 35940(sec)
2007/09/23: Pumping on going, No Observation / night time = 35760(sec)
2007/09/24: / night time = 0(sec)
2007/09/25: / night time = 0(sec)
2007/09/26: Dark 2733, 2734, 2735 ok / night time = 34920(sec)
2007/09/27: D 2736 (60s) OK, D2737 (300s) OK / night time = 34920(sec)
2007/09/28: / night time = 0(sec)
2007/09/29: / night time = 0(sec)
2007/09/30: Took a while to setup after power failure. Bad seeing. D2738 (60 sec) / night time = 34320(sec)
TOTAL NIGHT TIME: 974220(sec)
OPERATING RATE: 53.3 %
GB: 32.0 %
LMC: 20.5 %
SMC: 0.8 %
ALERT: 0.0 %
GRB: 0.0 %
Return to index
Previous month
Next month